# Interpretable Graph Neural Networks for Tabular Data

### v2

@@ 1 Introduction
In some application domains, e.g., medicine and law, predictions made by machine learning models need justification for legal and ethical considerations (Lakkaraju et al. 2017; Goodman and Flaxman 2017). In addition, users may put trust in such models only with a proper understanding of the reasoning behind the predictions. A direct solution is to use learning algorithms that produce interpretable models, such as logistic regression (Berkson 1944), which provides both local (instance-specific) and global (model-level) explanations for the predictions. However, such algorithms often result in a substantial loss in predictive performance compared to algorithms that generate black-box models, e.g., XGBoost (Chen and Guestrin 2016), Random Forests (Breiman 2001), and deep learning algorithms (Pintelas, Livieris, and Pintelas 2020; Mori and Uchihira 2019). Post-hoc explanation techniques, e.g., SHAP (Lundberg and Lee 2017), LIME (Ribeiro, Singh, and Guestrin 2016), and Anchors (Ribeiro, Singh, and Guestrin 2018), have been put forward as tools to explain predictions of the black-box models. However, the explanations provided by such techniques are limited in that they either do not show how exactly the predictions are computed, but merely present feature scores, such as LIME and SHAP, or come with no guarantees on the fidelity, i.e., that the provided explanation agrees with the underlying model (Yeh et al. 2019; Delaunay, Galárraga, and Largouët 2020). As extensively argued in (Rudin 2019), there are hence several reasons to consider generating interpretable models in the first place, if trustworthiness is a central concern.
@@
Graph Neural Networks (GNNs) have emerged as a powerful framework for representation learning of graph-structured data (Xu et al. 2019). The application of GNNs has been extended to tabular data, where a GNN can be used to learn an enhanced representation for the data points (rows) or to model the interaction between different features (columns). TabGNN (Guo et al. 2021) is an example of the first approach, where each data point is represented as a node in a graph. In comparison, TabularNet (Du et al. 2021) and Table2Graph (Zhou et al. 2022) follow the second approach, where the first uses a Graph Convolutional Network to model the relationships between features, and the second learns a probability adjacency matrix for a unified graph that models the interaction between features of the data points. GNNs can also be combined with other algorithms suited for tabular data, e.g., as in BGNN (Ivanov and Prokhorenkova 2021), which combines gradient-boosted decision trees and a GNN in one pipeline, where the GNN addresses the graph structure and the gradient-boosted decision trees handle the heterogeneous features of the tabular data. To the best of our knowledge, all previous approaches to using GNNs for tabular data result in black-box models and they are hence associated with the issues discussed above when applied in contexts with strong requirements on trustworthiness. In this work, we propose a novel GNN approach for tabular data, with the aim to eliminate the need to apply post-hoc explanation techniques without sacrificing predictive performance.
@@
  • a large-scale empirical investigation evaluating the explanations of IGNNet as well as comparing the predictive performance of IGNNet to state-of-the-art approaches for tabular data; XGBoost, Random Forests, and multi-layer perceptron (MLP), as well as to an algorithm generating interpretable models; TabNet (Arik and Pfister 2021)
@@ 2 Related Work
In this section, we provide some pointers to self-explaining (regular and graph) neural networks, and briefly discuss their relation to model-agnostic explanation techniques and interpretable models. We also provide pointers to work on interpretable deep learning approaches for tabular data.
@@ 2.1 Self-Explaining Neural Networks
Approaches in this category include the method in (Lei, Barzilay, and Jaakkola 2016), which is an early self-explaining neural network for text classification, the Contextual Explanation Network (CEN) (Al-Shedivat, Dubey, and Xing 2022), which generates explanations using intermediate graphical models, the Self-explaining Neural Network (SENN) (Alvarez Melis and Jaakkola 2018), which generalizes linear classifiers to neural networks using a concept autoencoder, and the CBM-AUC (Sawada and Nakamura 2022), which improves the efficiency of the former by replacing the decoder with a discriminator. Some approaches generate explanations in the form of counterfactual examples, e.g., CounterNet (Guo, Nguyen, and Yadav 2021), and Variational Counter Net (VCNet) (Guyomard, Fessant, and Guyet 2022). Again, such explanations do not provide detailed information on how the original predictions are computed and how exactly the input features affect the outcome.
@@ 2.2 Self-Explaining Graph Neural Networks
The Self-Explaining GNN (SE-GNN) (Dai and Wang 2021) uses similarities between nodes to make predictions on the nodes’ labels and provide explanations using the most similar K nodes with labels. ProtGNN (Zhang et al. 2022) also computes similarities, but between the input graph and prototypical graph patterns that are learned per class. Cui et al. (Cui et al. 2022) proposed a framework to build interpretable GNNs for connectome-based brain disorder analysis that resembles the signal correlation between different brain areas.
Similar to the (standard) self-explaining neural networks, the approaches that target graph neural networks provide abstract views of the predictions; the users cannot trace the exact computations, in contrast to when using interpretable models, which in principle allows for the inferences to be executed by hand.
@@ 2.3 Interpretable Deep Learning for Tabular Data
In an endeavor to provide an interpretable regression model for tabular data while retaining the performance of deep learning models, and inspired by generalized linear models (GLM), LocalGLMnet was proposed to make the regression parameters of a GLM feature dependent, allowing for quantifying variable importance and also conducting variable selection (Richman and Wüthrich 2022). TabNet (Arik and Pfister 2021) is another interpretable method proposed for tabular data learning, which employs a sequential attention mechanism and learnable masks for selecting a subset of meaningful features to reason from at each decision step. The feature selection is instance-based, i.e., it differs from one instance to another. The feature selection masks can be visualized to highlight important features and show how they are combined. However, it is not obvious how the features are actually used to form the predictions.
@@ 3.1 Interpretable Graph Neural Network
The input to a GNN learning algorithm is a set of graphs denoted by $\mathcal{G}=(V,E,X,\mathcal{W})$, consisting of a set of nodes $V$, a set of edges $E$, a set of node feature vectors $X$, and a set of edge weights $\mathcal{W}$, where $V=\{v_{1},\ldots,v_{N}\}$, $E\subseteq\{(v_{i},v_{j})|v_{i},v_{j}\in V\}$, $X=\{\textbf{x}_{1},\ldots,\textbf{x}_{N}\}$, and the weight of edge $(v_{i},v_{j})$ is represented by a scalar value $\delta_{i,j}$ in the set of edge weights $\mathcal{W}$, where $\delta_{i,j}=\mathcal{W}(i,j)$. A GNN algorithm learns a representation vector $\textbf{h}_{i}$ for each node $v_{i}$, which is initialized as $\textbf{h}_{i}^{(0)}=\textbf{x}_{i}$. The key steps in a GNN for graph classification can be summarized by the following two phases (Xu et al. 2019):
@@
    $\textbf{h}_{i}^{(l+1)}=\varphi\left(\textbf{w}^{(l+1)}\left(\delta_{i,i}\textbf{h}_{i}^{(l)}+\sum_{u\in\mathcal{N}(i)}\delta_{i,u}\textbf{h}_{u}^{(l)}\right)\right)$ (1)
    where $\delta_{i,u}$ is the weight assigned to the edge between node $v_{i}$ and node $v_{u}$. $\textbf{h}_{i}^{(l)}$ is the hidden representation of the node $v_{i}$ in the $l$-th layer, $\textbf{w}^{(l+1)}$ represents the learnable parameters, and $\varphi$ is a non-linearity function.
    The adjacency matrix A of size $|V|\times|V|$ contains the edge weights and can be normalized similar to a Graph Convolutional Network (GCN) (Kipf and Welling 2017) as shown in (2).
@@
    Here D is the degree matrix $\textit{{D}}_{ii}=\sum_{j}\textit{{A}}_{ij}$ (Kipf and Welling 2017).
@@
    Graph Pooling (Readout): A representation of the whole graph $\mathcal{G}$ is learned using a simple or advanced function (Xu et al. 2019), e.g, sum, mean, or MLP.
The whole graph representation obtained from the graph pooling phase can be submitted to a classifier to predict the class of the graph, which can be trained in an end-to-end architecture (Zhang et al. 2018; Ying et al. 2018).
The pooling function can be designed to provide an interpretable graph classification layer. Thus, the final hidden representation of each node is mapped using an injective readout function to a single value, for instance, through a neural network layer or dot product ($\mathcal{R}(\textbf{h}_{i}^{(l+1)}\in\mathbb{R}^{n})=h_{i}\in\mathbb{R}^{1}$), and concatenated to obtain the final representation g of the graph $\mathcal{G}$ where a scalar value in g corresponds to a node in the graph. Consequently, if a set of weights is applied to classify the graph, we can trace the contribution of each node to the predicted outcome, i.e., the user can find out which nodes contributed to the predicted class. For example, g can be used directly as follows:
$\hat{y}=\text{link}\left(\sum_{i=1}^{n}{\textbf{w}_{i}\textbf{g}_{i}}\right)$ (3)
where $\textbf{w}_{i}$ is the weight assigned to node $v_{i}$ represented in $\textbf{g}_{i}$.The link function is applied to accommodate a valid range of outputs, e.g., the sigmoid function for binary and softmax for multi-class classification. This is equivalent to:
$\hat{y}=\text{link}\left(\sum_{i=1}^{n}{\textbf{w}_{i}\mathcal{R}(\textbf{h}_{i}^{(l+1)})}\right)$ (4)
@@ 3.2 Representing Tabular Data Points as Graphs
Data: a set of graphs $\mathbb{G}$ and labels $y$
@@
             $\mathcal{L}\leftarrow\mathcal{L}+$loss$(\hat{y}_{j},y_{j})$
@@ 3.3 How can IGNNet achieve high performance while maintaining interpretability?
An expressive GNN can potentially capture complex patterns and dependencies in the graph, allowing nodes to be mapped to distinct representations based on their characteristics and relationships (Li and Leskovec 2022). Moreover, a GNN with an injective aggregation scheme can not only distinguish different structures but also map similar structures to similar representations (Xu et al. 2019). Therefore, if the tabular data are properly presented as graphs, GNNs with the aforementioned expressive capacities can model relationships and interactions between features, and consequently approximate complex non-linear mappings from inputs to predictions. On top of that, it has been shown by (ENNADIR et al. 2023) that GCNs based on 1-Lipschitz continuous activation functions can be improved in stability and robustness with Lipschitz normalization and continuity analysis; similar findings have also been demonstrated on graph attention networks (GAT) (Dasoulas, Scaman, and Virmaux 2021). This property is of particular importance when the application domain endures adversarial attacks or incomplete tabular data.
@@
    Each node is bounded to interact with a particular neighborhood, where it maintains correlations with the nodes within that neighborhood. As a result, the aggregated messages could potentially hold significance to the input feature values.
Since the edge weights are the correlation values, they determine the strength and the sign of the messages obtained from a neighborhood, allowing each node to store information not only concerning the original input feature but also features that are correlated with it. The proposed graph pooling function, combined with the constrained message-passing layers that keep representative information about the input features, allows tracking each feature’s contribution at the output layer and also through the message-passing layers all the way to the input features.
@@ 4.1 Experimental Setup
A GNN consists of one or more message-passing layers, and each layer can have a different design, e.g., different activation functions and batch normalization, which is the intra-layer design level (You, Ying, and Leskovec 2020). There is also the inter-layer design level which involves, for instance, how the message-passing layers are organized into a neural network and if the skip connections are added between layers (You, Ying, and Leskovec 2020). While the intra-layer and inter-layer designs can vary based on the nature of the prediction task, we propose a general architecture for our empirical investigation. [^1] The source code is available here: https://github.com/amrmalkhatib/ignnet However, it is up to the user to modify the architecture of the GNN. The number of message-passing layers, number of units in linear transformations, and other hyperparameters were found based on a quasi-random search and evaluation on development sets of the following three datasets: Churn, Electricity, and Higgs. We have six message-passing layers in the proposed architecture, each with a Relu activation function. Multiple learnable weights are also applied to the nodes’ representation, followed by a Relu function. Besides three batch normalization layers, four skip connections are added as illustrated in Figure 2. After all the GNN layers, we use a feedforward neural network (FNN) to map the multidimensional representation of each node into a single value. In the FNN, we do not include any activation functions in order to keep the mapping linear, but a sigmoid function is applied after the final layer to obtain a value between 0 and 1 for each node. The FNN is composed of 8 layers with the following numbers of units (128, 64, 32, 16, 8, 4, 2, 1) and 3 batch normalization layers after the second, fourth, and sixth hidden layers. After the FNN, the nodes’ final values are concatenated to form a representation of the whole graph (data instance). Finally, the weights that are output are used to make predictions. The GNN is trained end-to-end, starting from the embeddings layer and ending with the class prediction.
Figure 2: IGNNet default architecture. It starts with the embedding layer, a linear transformation from one dimension to 64 dimensions. A Relu activation function follows each message-passing layer and each green block as well. The feedforward network at the end has no activation functions between layers to ensure a linear transformation into a single value. A sigmoid activation function follows the feedforward network to obtain the final value for each feature between 0 and 1.
@@
In the experiments, 35 publicly available datasets are used. [^2] All the datasets were obtained from https://www.openml.org Each dataset is split into training, development, and test sets. The development set is used for overfitting detection and early stopping of the training process, the training set is used to train the model, and the test set is used to evaluate the model. [^3] Detailed information about each dataset is provided in the appendix. For a fair comparison, all the compared learning algorithms are trained without hyperparameters tuning using the default settings on each dataset. In cases where the learning algorithm does not employ the development set to measure performance progress for early stopping, the development and training subsets are combined into a joint training set. The adjacency matrix uses the correlation values computed on the training data split. The weight on edge from the node to itself (self-loop) is a user-adjustable hyperparameter, constitutes between 70% to 90% (on average) of the weighted summation to keep a strong message per node that does not fade out with multiple layers of message-passing. Weak correlation values are excluded from the graph, so if the absolute correlation value is below 0.2, the edge is removed unless no correlation values are above 0.2; in case of the latter, the procedure is repeated using a reduced threshold of 0.05.444In the technical appendix, we provide an ablation study on the effects of different hyperparameter preferences on predictive performance. The Pearson correlation coefficient (Pearson 1895) is used to estimate the linear relationship between features. In the data preprocessing step, the categorical features are binarized using one-hot encoding, and all the feature values are normalized using min-max normalization (the max and min values are computed on the training split). The normalization keeps the feature values between 0 and 1, which is essential for the IGNNet to have one scale for all nodes.
@@
For imbalanced binary classification datasets, we oversample the minority class in the training set to align the size with the majority class. All the compared algorithms are trained using the oversampled training data. While for multi-class datasets, no oversampling is conducted. The area under the ROC curve (AUC) is used to measure the predictive performance, as it is not affected by the decision threshold. For the multi-class datasets, weighted AUC is calculated, i.e., the AUC is computed for each class against the rest and weighted by the support.

### v3

@@ 1 Introduction
In some application domains, e.g., medicine and law, predictions made by machine learning models need justification for legal and ethical considerations [24, 17]. In addition, users may put trust in such models only with a proper understanding of the reasoning behind the predictions. A direct solution is to use learning algorithms that produce interpretable models, such as logistic regression [5], which provides both local (instance-specific) and global (model-level) explanations for the predictions. However, such algorithms often result in a substantial loss in predictive performance compared to algorithms that generate black-box models, e.g., XGBoost [7], Random Forests [6], and deep learning algorithms [29, 32]. Post-hoc explanation techniques, e.g., SHAP [27], LIME [34], and Anchors [35], have been put forward as tools to explain predictions of the black-box models. However, the explanations provided by such methods are often computationally intensive [2, 22] and lack fidelity guarantees, i.e., there are no guarantees that the provided explanation accurately reflects the underlying model, as studied previously [8, 12, 42]. As extensively argued in [37], there are hence several reasons to consider generating interpretable models in the first place, if trustworthiness is a central concern.
@@
Graph Neural Networks (GNNs) have emerged as a powerful framework for representation learning of graph-structured data [40]. The application of GNNs has been extended to tabular data, where a GNN can be used to learn an enhanced representation for the data points (rows) or to model the interaction between different features (columns). TabGNN [19] is an example of the first approach, where each data point is represented as a node in a graph. In comparison, TabularNet [13] and Table2Graph [47] follow the second approach, where the first uses a Graph Convolutional Network to model the relationships between features, and the second learns a probability adjacency matrix for a unified graph that models the interaction between features of the data points. GNNs can also be combined with other algorithms suited for tabular data, e.g., as in BGNN [21], which combines gradient-boosted decision trees and a GNN in one pipeline, where the GNN addresses the graph structure and the gradient-boosted decision trees handle the heterogeneous features of the tabular data. To the best of our knowledge, all previous approaches to using GNNs for tabular data result in black-box models and they are hence associated with the issues discussed above when applied in contexts with strong requirements on trustworthiness. In this work, we propose a novel GNN approach for tabular data, with the aim to eliminate the need to apply post-hoc explanation techniques without sacrificing predictive performance.
@@
  • a large-scale empirical investigation evaluating the explanations of IGNNet as well as comparing the predictive performance of IGNNet to state-of-the-art approaches for tabular data; XGBoost, Random Forests, and multi-layer perceptron (MLP), as well as to an algorithm generating interpretable models; TabNet [4]
@@ 2 Related Work
In this section, we provide pointers to self-explaining (regular and graph) neural networks, and briefly discuss their relation to model-agnostic explanation techniques and interpretable models. We also provide pointers to work on interpretable deep learning approaches for tabular data.
@@ 2.1 Self-Explaining Neural Networks
Approaches in this category include the method in [25], which is an early self-explaining neural network for text classification, the Contextual Explanation Network (CEN) [1], which generates explanations using intermediate graphical models, the Self-explaining Neural Network (SENN) [3], which generalizes linear classifiers to neural networks using a concept autoencoder, and the CBM-AUC [38], which improves the efficiency of the former by replacing the decoder with a discriminator. Some approaches generate explanations in the form of counterfactual examples, e.g., CounterNet [18], and Variational Counter Net (VCNet) [20]. Again, such explanations do not provide detailed information on how the original predictions are computed and how exactly the input features affect the outcome.
@@ 2.2 Self-Explaining Graph Neural Networks
The Self-Explaining GNN (SE-GNN) [10] uses similarities between nodes to make predictions on the nodes’ labels and provide explanations using the most similar K nodes with labels. ProtGNN [46] also computes similarities, but between the input graph and prototypical graph patterns that are learned per class. Cui et al. [9] proposed a framework to build interpretable GNNs for connectome-based brain disorder analysis that resembles the signal correlation between different brain areas. Feng et al. [15] proposed the KerGNN (Kernel Graph Neural Network), which improves model interpretability using graph filters. The graph filters learned in KerGNNs can uncover local graph structures in a dataset. Xuanyuan et al. [41] proposed scrutinizing individual neurons in a GNN to generate global explanations using neuron-level concepts. CLARUS [28] enhances human understanding of GNN predictions in medicine and facilitates the visualization of patient-specific information.
The self-explainable GNN methods mentioned above are not suited for tabular data, as they are designed for input data that are inherently graphical, such as social networks and molecular structures.
@@ 2.3 Interpretable Deep Learning for Tabular Data
In an endeavor to provide an interpretable regression model for tabular data while retaining the performance of deep learning models, and inspired by generalized linear models (GLM), LocalGLMnet was proposed to make the regression parameters of a GLM feature dependent, allowing for quantifying variable importance and also conducting variable selection [36]. TabNet [4] is another interpretable method proposed for tabular data learning, which employs a sequential attention mechanism and learnable masks for selecting a subset of meaningful features to reason from at each decision step. The feature selection is instance-based, i.e., it differs from one instance to another. The feature selection masks can be visualized to highlight important features and show how they are combined. However, it is not obvious how the features are actually used to form the predictions.
@@ 3.1 Interpretable Graph Neural Network
The input to a GNN learning algorithm is a set of graphs denoted by $\mathcal{G}=(V,E,X,\textit{{A}})$, consisting of a set of nodes $V$, a set of edges $E$, a set of node feature vectors $X$, and an adjacency matrix A, where $V=\{v_{1},\ldots,v_{N}\}$, $E\subseteq\{(v_{i},v_{j})|v_{i},v_{j}\in V\}$, $X=\{\textbf{x}_{1},\ldots,\textbf{x}_{N}\}$, and the weight of edge $(v_{i},v_{j})$ is represented by a scalar value $\delta_{i,j}$ in the weighted adjacency matrix A, where $\delta_{i,j}=\textit{{A}}(i,j)$. A GNN algorithm learns a representation vector $\textbf{h}_{i}$ for each node $v_{i}$, which is initialized as $\textbf{h}_{i}^{(0)}=\textbf{x}_{i}$. The key steps in a GNN for graph classification can be summarized by the following two phases [40]:
@@
    $\textbf{h}_{i}^{(l+1)}=\varphi\left(\textbf{w}^{(l)}\left(\delta_{i,i}\textbf{h}_{i}^{(l)}+\sum_{u\in\mathcal{N}(i)}\delta_{i,u}\textbf{h}_{u}^{(l)}\right)\right)$ (1)
    Where $\textbf{h}_{i}^{(l)}$ is the hidden representation of the node $v_{i}$ in the $l$-th layer, $\mathcal{N}(i)$ is the neighborhood of node $v_{i}$, $\textbf{w}^{(l)}$ represents the learnable parameters, and $\varphi$ is a non-linearity function.
    The adjacency matrix A of size $|V|\times|V|$ contains the edge weights and can be normalized similar to a Graph Convolutional Network (GCN) [23] as shown in (2).
@@
    Here D is the degree matrix $\textit{{D}}_{ii}=\sum_{j}\textit{{A}}_{ij}$ [23].
@@
    Graph Pooling (Readout): A representation of the whole graph $\mathcal{G}$ is learned using a simple or advanced function [40], e.g, sum, mean, or MLP.
The whole graph representation obtained from the graph pooling phase can be submitted to a classifier to predict the class of the graph, which can be trained in an end-to-end architecture [45, 43].
The pooling function can be designed to provide an interpretable graph classification layer. Thus, the final hidden representation of each node is mapped to a single value, for instance, through a neural network layer or dot product ($\mathcal{R}(\textbf{h}_{i}^{(l+1)}\in\mathbb{R}^{n})=h_{i}\in\mathbb{R}^{1}$), and concatenated to obtain the final representation g of the graph $\mathcal{G}$ where a scalar value in g corresponds to a node in the graph. Consequently, if a set of weights is applied to classify the graph, we can trace the contribution of each node to the predicted outcome, i.e., the user can find out which nodes contributed to the predicted class. For example, g can be used directly as follows:
$\hat{y}=\text{link}\left(\sum_{i=1}^{n}{w_{i}g_{i}}\right)$ (3)
where $w_{i}$ is the weight (in vector w) assigned to node $v_{i}$ represented in $g_{i}$. The link function is applied to accommodate a valid range of outputs, e.g., the sigmoid function for binary and softmax for multi-class classification. This is equivalent to:
$\hat{y}=\text{link}\left(\sum_{i=1}^{n}{w_{i}\mathcal{R}(\textbf{h}_{i}^{(l+1)})}\right)$ (4)
@@ 3.2 Representing Tabular Data Points as Graphs
Data: a set of graphs $\mathbb{G}$ and labels $\mathbb{Y}$
@@
             $\mathcal{L}\leftarrow\mathcal{L}+$loss$(\hat{y}_{j},y_{j}\in\mathbb{Y})$
@@ 3.3 How can IGNNet achieve high performance while maintaining interpretability?
An expressive GNN can potentially capture complex patterns and dependencies in the graph, allowing nodes to be mapped to distinct representations based on their characteristics and relationships [26]. Moreover, a GNN with an injective aggregation scheme can not only distinguish different structures but also map similar structures to similar representations [40]. Therefore, if the tabular data are properly presented as graphs, GNNs with the aforementioned expressive capacities can model relationships and interactions between features, and consequently approximate complex non-linear mappings from inputs to predictions. On top of that, it has been shown by [14] that GCNs based on 1-Lipschitz continuous activation functions can be improved in stability and robustness with Lipschitz normalization and continuity analysis; similar findings have also been demonstrated on graph attention networks (GAT) [11]. This property is of particular importance when the application domain endures adversarial attacks or incomplete tabular data.
@@
    Each node has a highly weighted self-loop which conveys the main message of the node.
  3. 3.
    Each node is bounded to interact with a particular neighborhood, where it maintains correlations with the nodes within that neighborhood.
The message-passing operation is based on the linear relationships between nodes. The strength of a passed message is directly proportional to the linear relationship between the two nodes, and the sign of the correlation value determines the signs of the messages. In addition, the highly weighted self-loops maintain the main messages carried by the nodes and prevent them from fading away through multiple message-passing layers. As a result, the aggregated messages could potentially hold significance to the input feature values. The proposed graph pooling function, combined with the constrained message-passing layers that keep representative information about the input features, allows tracking each feature’s contribution at the output layer and also through the message-passing layers all the way to the input features.
@@ 4.1 Experimental Setup
A GNN consists of one or more message-passing layers, and each layer can have a different design, e.g., different activation functions and batch normalization, which is the intra-layer design level [44]. There is also the inter-layer design level, which involves, for instance, how the message-passing layers are organized into a neural network and if skip connections are added between layers [44]. While the intra-layer and inter-layer designs can vary based on the nature of the prediction task, we propose a general architecture for our empirical investigation. [^1] The source code is available here: https://github.com/amrmalkhatib/ignnet However, it is up to the user to modify the architecture of the GNN. The number of message-passing layers, number of units in linear transformations, and other hyperparameters were found based on a quasi-random search and evaluation on development sets of the following three datasets: Churn, Electricity, and Higgs. We have six message-passing layers in the proposed architecture, each with a Relu activation function. Multiple learnable weights are also applied to the nodes’ representation, followed by a Relu function. Besides three batch normalization layers, four skip connections are added as illustrated in Figure 2. After all the GNN layers, we use a feedforward neural network (FNN) to map the multidimensional representation of each node into a single value. In the FNN, we do not include any activation functions in order to keep the mapping linear, but a sigmoid function is applied after the final layer to obtain a value between 0 and 1 for each node. The FNN is composed of 8 layers with the following numbers of units (128, 64, 32, 16, 8, 4, 2, 1) and 3 batch normalization layers after the second, fourth, and sixth hidden layers. After the FNN, the nodes’ final values are concatenated to form a representation of the whole graph (data instance). Finally, the weights that are output are used to make predictions. The GNN is trained end-to-end, starting from the embeddings layer and ending with the class prediction.
Figure 2: IGNNet default architecture. It starts with a layer to project the features into higher dimensionality, a linear transformation from one dimension to 64 dimensions. A Relu activation function follows each message-passing layer and each green block as well. The feedforward network at the end has no activation functions between layers to ensure a linear transformation into a single value. A sigmoid activation function follows the feedforward network to obtain the final value for each feature between 0 and 1.
@@
In the experiments, 35 publicly available datasets are used. [^2] All the datasets were obtained from https://openml.org Each dataset is split into training, development, and test sets. The development set is used for overfitting detection and early stopping of the training process, the training set is used to train the model, and the test set is used to evaluate the model. [^3] Detailed information about each dataset is provided in the appendix For a fair comparison, all the compared learning algorithms are trained without hyperparameters tuning using the default settings on each dataset. In cases where the learning algorithm does not employ the development set to measure performance progress for early stopping, the development and training subsets are combined into a joint training set. The adjacency matrix uses the correlation values computed on the training data split. The weight on edge from the node to itself (self-loop) is a user-adjustable hyperparameter, constitutes between 70% to 90% (on average) of the weighted summation to keep a strong message per node that does not fade out with multiple layers of message-passing. Weak correlation values are excluded from the graph, so if the absolute correlation value is below 0.2, the edge is removed unless no correlation values are above 0.2; in case of the latter, the procedure is repeated using a reduced threshold of 0.05.444The appendix includes an ablation study on how various hyperparameter preferences affect predictive performance. The Pearson correlation coefficient [31] is used to estimate the linear relationship between features. In the data preprocessing step, the categorical features are binarized using one-hot encoding, and all the feature values are normalized using min-max normalization (the max and min values are computed on the training split). The normalization keeps the feature values between 0 and 1, which is essential for the IGNNet to have one scale for all nodes.
@@
For imbalanced binary classification datasets, we randomly oversample the minority class in the training set to align the size with the majority class. All the compared algorithms are trained using the oversampled training data. While for multi-class datasets, no oversampling is conducted.
The area under the ROC curve (AUC) is used to measure the predictive performance, as it is not affected by the decision threshold. For the multi-class datasets, weighted AUC is calculated, i.e., the AUC is computed for each class against the rest and weighted by the support.